### v2

# Discrete-Time CRLB-based Power Allocation for CF MIMO-ISAC with Joint Localization and Velocity Sensing

Abstract
In this paper, we investigate integrated sensing and communication (ISAC) in a cell-free (CF) multiple-input multiple-output (MIMO) network, where each access point functions either as an ISAC transmitter or as a sensing receiver. We devote into the ISAC sensing metric using the discrete-time signal-based Cramér–Rao lower bounds (CRLBs) for joint location and velocity estimation under arbitrary power allocation ratios under the deterministic radar cross section assumption (RCS). Then, we consider the power allocation optimization problem for the CF MIMO-ISAC as the maximization of the communication signal-to-interference-plus-noise ratio (SINR), subject to CRLB-based sensing constraints and per-transmitter power limits. To solve the resulting nonlinear and non-convex problem, we propose a penalty function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS), and an alternative method based on a modified steepest descent approach (PP-MSD-ILS). We show that the proposed algorithms are scalable and can be extended to a broad class of optimization problems involving nonlinear inequality constraints and affine equality constraints. In addition, we extend the PP-MCG-ILS algorithm to the pure sensing scenario, where a penalty function-based normalized conjugate gradient algorithm (P-NCG-ILS) is developed for sensing power minimization. Finally, we analyze the convergence behavior and qualitatively compare the computational complexity of the proposed algorithms. Simulation results confirm the accuracy of the derived CRLBs and demonstrate the effectiveness of the proposed power allocation strategies in enhancing both sensing and overall ISAC performance.
Index Terms:
Cell-free MIMO ISAC, CRLB, Power allocation, Penalty function, Projection operator, Conjugate gradient.
I Introduction
Integrated sensing and communication (ISAC) has emerged as a key enabling technology for sixth-generation (6G) applications, including smart cities, autonomous driving, and remote healthcare [1]. Unlike traditional systems that separate the design of communication and sensing functionalities, ISAC integrates both within a unified framework using shared hardware platforms and spectral resources. This integration offers substantial improvements in spectral efficiency, energy savings, and hardware utilization [2, 3].
ISAC systems can be broadly categorized based on the spatial configuration of transmitters and receivers. Monostatic ISAC systems, where the transmitter and receiver are co-located, typically rely on advanced beamforming techniques to mitigate self-interference in full-duplex operations [4, 5]. In contrast, multistatic ISAC systems deploy spatially separated transmitters and receivers, facilitating spatial diversity and sensing coverage [6, 7]. Building upon the multistatic architecture, the concept of cell-free (CF) multiple-input and multiple-output (MIMO) ISAC has recently emerged [8, 9]. In CF MIMO-ISAC systems, a large number of distributed access points (APs) jointly serve communication users and provide sensing services in a coordinated manner without relying on conventional cell boundaries. This architecture inherits the advantages of both distributed multistatic radar and CF MIMO communication, offering improved spectral efficiency, spatial resolution, and system scalability [10].
Recent advancements in CF MIMO-ISAC have extended from basic trade-off analysis [11] to complex sensing tasks, including target detection [12], and parameter estimation for radar cross-section (RCS) [13], location [14], velocity [15], and angle [16]. For example, [12] studies the fundamental performance tradeoff between the target detection probability and the user’s achievable rate in ISAC systems by deriving the probability of false alarm and the successful probability of detection for the target of interest. [13] provides a unified framework integrating sensing (RCS estimation), computing, and communication to optimize limited system resource. [14] proposes a two-stage scheme for target localization, with communication signals reused as sensing reference signals. Additionally, joint sensing and communication beamforming has been explored in [17].
Despite these advances, a fundamental challenge in ISAC system design, including CF MIMO-ISAC, is the formulation of robust performance metrics, which are critical for guiding system design and balancing communication-sensing trade-offs [2, 3]. While unified metrics based on mutual information [18] and Kullback–Leibler divergence [19] have been proposed for general analysis, adapting these metrics to task-specific sensing objectives remains a significant challenge. Consequently, task-specific metrics for sensing and communication are still widely adopted.
For communication, Shannon capacity underpins metrics such as spectral efficiency [20], achievable rate [21], and signal-to-interference-plus-noise ratio (SINR). Additional metrics, including coverage probability and energy efficiency, have also been considered, particularly in scenarios with stringent quality-of-service (QoS) or energy constraints [22]. On the sensing side, a variety of performance metrics have been investigated, such as sensing mutual information [23], sensing SINR [24, 25], and the Cramér–Rao lower bound (CRLB) [26, 27, 28, 29]. Among these, CRLB-based formulations have proven particularly effective for parameter estimation and algorithm benchmarking. CRLB-based approaches generally fall into two categories depending on the RCS model: deterministic and stochastic. Under the stochastic RCS assumption, RCS values are modeled as random variables, and the corresponding likelihood functions are derived by averaging over their statistical distributions. In contrast, the deterministic RCS assumption treats RCS values as deterministic but unknown quantities that may vary spatially. This model is particularly applicable in slowly varying environments or when prior RCS knowledge is unavailable. CRLBs for both single-target and multiple-target location estimation have been analyzed under the stochastic RCS assumption [26, 27] and the deterministic RCS assumption [28, 29].
However, most existing CRLB formulations are built upon continuous-time signal models, whereas practical radar systems operate in the discrete-time domain, where the sampling rate critically affects sensing accuracy. Our prior work [30] addresses this gap by introducing a discrete-time likelihood function over Gaussian channels, offering a more realistic modeling framework. Building on this foundation, this paper devotes into the CRLBs derivation for joint location and velocity estimation under the deterministic RCS assumption, covering both single- and multi-target cases in CF MIMO-ISAC systems.
Beyond performance characterization, another critical challenge lies in resource allocation, involving time [31], frequency [32], space [33], and power [24, 25, 34, 35, 36]. Recent works in CF MIMO-ISAC systems have addressed power allocation (PA) under sensing SINR constraints [24, 25] or sensing CRLB constraints [34, 35, 36]. These involved sensing CRLBs typically consider localization only, neglecting velocity or angular estimation, and depend on continuous-time models that overlook the impact of sampling rate.
These considerations motivate us to consider a more comprehensive and realistic setting: joint location and velocity estimation in CF MIMO-ISAC systems under a discrete-time signal model with a deterministic RCS assumption. The CRLBs are serve as effective sensing metrics to guide PA across distributed APs, with the goal of maximizing communication SINR. This leads to a class of non-convex and highly nonlinear optimization problems. To address this issue, we develop a tailored optimization framework that integrates penalty methods, projection operations, and gradient-based search, enabling efficient and reliable algorithmic solutions under complex practical constraints. The main contributions are as follows:
  • We derive novel closed-form approximate CRLBs for joint location and velocity estimation under arbitrary PA strategies. These CRLBs are rigorously validated against maximum likelihood estimation (MLE) results, confirming their effectiveness as practical sensing performance metrics. Based on these metrics, we formulate the PA optimization problems for ISAC, aiming at communication SINR maximization while ensuring joint CRLB and power budget constraints.
  • To tackle the non-convexity and nonlinearity in PA optimization for ISAC, we propose a penalty-function and projection-based modified conjugate gradient algorithm with inexact line search (PP-MCG-ILS) together with PP-MSD-ILS algorithm based on the modified steepest descent algorithm (MSD). We further propose a penalty-function-based normalized modified conjugate gradient algorithm with inexact line search (P-NCG-ILS), which can effectively solve the power minimization problem subject to sensing accuracy constraints.
  • We provide a comprehensive analysis of their convergence behavior and computational complexity. The design ensures stable updates, rapid convergence and low computational overhead. Additionally, we show that the proposed PP-MCG-ILS and P-NCG-ILS algorithms can be broadly scalable to a class of constrained non-convex optimization problems involving affine equality constraints (handled via projection) and nonlinear inequality constraints (handled via penalties).
  • Extensive simulations validate the performance of the developed algorithms in both ISAC and sensing-only scenarios. The PP-MCG-ILS algorithm consistently exhibits fast convergence, robust numerical stability, and optimal PA efficiency across various conditions, as compared to benchmark methods. Results also highlight important trade-offs between sensing and communication performance and confirm that the proposed methods can meet stringent sensing QoS while maximizing communication throughput or minimizing power usage.
Section II introduces the received signal model for the CF MIMO-ISAC system. Section III-B formulates the power allocation problem for CF MIMO-ISAC. Section IV presents the proposed power allocation algorithms for both CF MIMO-ISAC and the sensing-only scenario. Section VI analyzes the convergence properties and computational complexity of the proposed algorithms. Section VII provides simulation results to validate the theoretical analysis and algorithmic performance. Finally, Section VIII concludes the paper.
[…]
II Signal Model
As shown in Fig. 1, we consider a CF MIMO-ISAC system with widely separated APs. Each AP either serves as an ISAC transmitter or a sensing receiver, similar to the setup in [24]. Assume that there are $N$ single-antenna transmitters and $K$ single-antenna receivers, with the positions of the transmitter $n$ and receiver $k$ denoted as ${\bm{l}}_{n}=[x_{n},y_{n}]^{\rm T}$ and ${\bm{l}}_{k}=[x_{k},y_{k}]^{\rm T}$, respectively. All transmitters and receivers are scheduled and managed by the edge cloud and the core network, e.g., the fifth generation (5G) communication networks. Denote the location and velocity vectors of the sensing target as ${\bm{l}}=[x,y]^{\rm T}$ and ${\bm{v}}=[v_{x},v_{y}]^{\rm T}$, respectively.
Figure 1: CF MIMO system for ISAC.
The Doppler frequency induced by the relative radial velocity of the target with respect to transmitter $n$ and receiver $k$ is computed as
$f_{n,k}=\dfrac{1}{\lambda}{\bm{v}}^{\rm T}\left(\frac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\frac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (1)
where $\lambda$ denotes the wavelength. Denote $\tau_{n,k}$ as the propagation delay between the $n$th transmitter and the $k$th receiver, as reflected by the target. Then, we have
$\tau_{n,k}$ $=\tau_{n}+\tau_{k},$ (2a)
$\tau_{n}={\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}/{c},$ $\quad\tau_{k}={\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}/{c},$ (2b)
where $\tau_{n}$ and $\tau_{k}$ represent the propagation delays from the $n$th transmitter to the $q$th target, and from the target to the $k$th receiver, respectively, and $c$ denotes the speed of light.
[…]
${\bm{x}}_{n}(t)=\sqrt{P\rho_{n}}{b}_{n}\varsigma s_{n}(t),$ (3)
where $P$ denotes the total available energy, $\rho_{n}$ is the PA factor with $\sum_{n=1}^{N}\rho_{n}=1$, ${b}_{n}\in\mathbb{C}$ is the normalized transmit beamforming weight with $|{b}_{n}|^{2}=1$, and $\varsigma\in\mathbb{C}$ has unit variance $\mathbb{E}\{|\varsigma|^{2}\}=1$. Let ${\bm{\rho}}=[\rho_{1},\rho_{2},\cdots,\rho_{N}]^{\rm T}$ be the vector of the PA coefficient of the $N$ transmitters.
Sensing model: For target-reflected reception, we refer the term “RCS” as encompassing both channel fading and radar RCS. Denote the RCS associated with transmitter $n$ and receiver $k$ as ${\alpha}_{n,k}$. We consider the Swerling-I sensing model [37], i.e. the deterministic RCS assumption where RCSs are assumed to be deterministic but unknown during the observation time [29, 28]. Then, the lowpass equivalent of the reflected signal from the $n$th transmitter by a target received at receiver $k$ is given by [^1] The transmitted signals over different transmitters are orthogonal and maintain approximate orthogonality over different delays and Doppler frequencies such that the signals contributed from different transmitters can be separated at each receiver [29, 26].
$r_{n,k}(t)=\alpha_{n,k}y_{n,k}(t)+z_{n,k}(t),$ (4)
where $z_{n,k}(t)$ denotes the clutter-plus-noise and $y_{n,k}(t)$ is the loss-free and noise-free received signal, i.e.,
${y}_{n,k}(t)=\sqrt{P\rho_{n}}b_{n}s_{n}(t-\tau_{n,k})e^{j2\pi f_{n,k}t}.$ (5)
Define the signal-to-clutter-plus-noise ratio (SCNR) as $\delta_{\rm cn}=\frac{P}{T_{\rm eff}\sigma_{\rm cn}^{2}}$ with $T_{\rm eff}$ denoting the effective waveform width.
Communication model: The received signal of the communication user is given by
$\small{r}_{u}(t)$ $=\sqrt{P}\sum\nolimits_{n=1}^{N}g_{n}{b}_{n}\sqrt{\rho_{n}}s_{n}(t)+{z}(t),$ (6)
where $g_{n}$ is the channel fading between the transmitter $n$ and the user, and $z(t)$ denotes the interference-plus-noise with zero mean and variance $\sigma_{z}^{2}$. Then, the SINR at the user terminal can be given by
$\small\gamma$ $=\sum\nolimits_{n=1}^{N}|g_{n}{b}_{n}|^{2}\rho_{n}\delta\overset{(\text{i})}{=}\delta{\bm{\rho}}^{\rm T}{\bm{g}}$ (7)
where $\delta=\frac{P}{T_{\rm eff}\sigma_{z}^{2}}$ is the total SINR and ${\bm{g}}$ is the channel gain [^2] The communication channels are assumed to be estimated a priori using pilot signals [24, 35, 34, 36]. with its $n$th element being $|g_{n}|^{2}$. Note that $(\mathrm{i})$ holds when using a normalized conjugate beamforming with ${b}_{n}={g}^{*}_{n}/|g_{n}|$.
III The Proposed Sensing Metric and Problem Formulation
In this section, we present an approximate CRLB for CF MIMO radar under arbitrary power allocations.
To begin with, we introduce the following definitions. Define the squared effective bandwidth (SEBW) and squared effective pulse time width (SETW) as
[…]
respectively.
III-A Sensing Metric
Unlike the sensing performance metrics employed in [24, 25, 34, 35, 36], which only consider target detection or location estimation, we consider the joint location and velocity CRLB as the sensing metric, developed based on the discrete-time reception of the reflected signal under the deterministic RCS assumption. In particular, we have derived the approximate CRLBs for both location and velocity in [30] for simplified sensing metrics. They are respectively given by
$\small{\bm{C}}_{\rm L}$ $=({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T})^{-1},$ (11)
and
$\small{\bm{C}}_{\rm V}$ $=({\bm{Y}}-{\bm{V}}^{\rm T}{\bm{P}}^{-1}{\bm{V}})^{-1}.$ (12)
The detailed expression of ${\bm{P}}$ and ${\bm{V}}$ are given at the top of the page, respectively, and
${\bm{P}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}^{2}a_{n,k}+2\beta_{n,k}\eta_{n,k}b_{n,k}+\eta_{n,k}^{2}d_{n,k}&\beta_{n,k}\zeta_{n,k}a_{n,k}\!+\!(\beta_{n,k}\kappa_{n,k}\!+\!\zeta_{n,k}\eta_{n,k})b_{n,k}+\eta_{n,k}\kappa_{n,k}d_{n,k}\\ \zeta_{n,k}\beta_{n,k}a_{n,k}\!+\!(\zeta_{n,k}\eta_{n,k}+\beta_{n,k}\kappa_{n,k})b_{n,k}+\kappa_{n,k}\eta_{n,k}d_{n,k}&\zeta_{n,k}^{2}a_{n,k}+2\zeta_{n,k}\kappa_{n,k}b_{n,k}\!+\!\kappa_{n,k}^{2}d_{n,k}\end{bmatrix},$
$\small{\bm{V}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\beta_{n,k}\xi_{n,k}b_{n,k}+\eta_{n,k}\xi_{n,k}d_{n,k}&\beta_{n,k}\varrho_{n,k}b_{n,k}+\eta_{n,k}\varrho_{n,k}d_{n,k}\\ \zeta_{n,k}\xi_{n,k}b_{n,k}+\kappa_{n,k}\xi_{n,k}d_{n,k}&\zeta_{n,k}\varrho_{n,k}b_{n,k}+\kappa_{n,k}\varrho_{n,k}d_{n,k}\end{bmatrix},$
$\small{\bm{Y}}=\sum_{n=1}^{N}\sum_{k=1}^{K}\begin{bmatrix}\xi_{n,k}^{2}d_{n,k}&\xi_{n,k}\varrho_{n,k}d_{n,k}\\ \varrho_{n,k}\xi_{n,k}d_{n,k}&\varrho_{n,k}^{2}d_{n,k}\end{bmatrix}.$ (13)
The relevant waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, and geometric-spreading-related factors $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, are detailed in Appendix A.
As observed from (61), the PA coefficients play a pivotal role in the formulation of the CRLBs. Meanwhile, these coefficients directly affect communication SINR, as described in (7). Unfortunately, the PA optimization relevant to the CRLB is intractable due to the implicit and complex forms of the expressions in (11) and (12). To facilitate analysis and optimization, we reformulate the CRLB expressions into a more explicit and tractable form.
We denote the $(i,j)$th elements of matrices ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$ as $p_{ij}$, $v_{ij}$ and $y_{ij}$, respectively. Then, one can simplify them as
$\small\begin{cases}p_{ij}=\sum_{n=1}^{N}w_{n}^{p_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{p_{ij}},\\ v_{ij}=\sum_{n=1}^{N}w_{n}^{v_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{v_{ij}},\\ y_{ij}=\sum_{n=1}^{N}w_{n}^{y_{ij}}\rho_{n}={\bm{\rho}}^{\rm T}{\bm{w}}_{y_{ij}},\end{cases}$ (14)
where $w_{n}^{p_{ij}}$, $w_{n}^{v_{ij}}$ and $w_{n}^{y_{ij}}$ are the $n$th element of the weight vector ${\bm{w}}_{p_{ij}}$, ${\bm{w}}_{v_{ij}}$ and ${\bm{w}}_{y_{ij}}$, respectively. These weights can be readily obtained from the definitions of ${\bm{P}}$, ${\bm{V}}$ and ${\bm{Y}}$, by extracting their PA coefficient $\rho_{n}$. For example, from (13), we have $w_{n}^{p_{11}}$ as
$\small w_{n}^{p_{11}}=\sum\nolimits_{k=1}^{K}(\beta_{n,k}^{2}\bar{a}_{n,k}+2\beta_{n,k}\eta_{n,k}\bar{b}_{n,k}+\eta_{n,k}^{2}\bar{d}_{n,k}),$ (15)
where the details of $\bar{a}_{n,k}$, $\bar{b}_{n,k}$ and $\bar{d}_{n,k}$ are referred to (61).
By substituting (14) into (11), the CRLB matrix for location estimation can be rewritten as
$\small{\bm{C}}_{\rm L}=\frac{1}{|{\bm{U}}|}\begin{bmatrix}u_{22}&-u_{12}\\ -u_{21}&u_{11}\end{bmatrix},$ (16)
where $|{\bm{U}}|=u_{11}u_{22}-u_{12}u_{21}$. The relevant elements $u_{ij}$ are given respectively by
$\small u_{11}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$ (17)
$u_{12}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{y}){\bm{\rho}}\big{]},$
$u_{22}=\frac{1}{|{\bm{Y}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{p_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{y}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{y}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{y}){\bm{\rho}}\big{]},$
where $|{\bm{Y}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{y}{\bm{\rho}}$, and
$\small\begin{cases}{\bm{W}}_{y}={\bm{w}}_{y_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{y_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{y_{21}}^{\rm T},\\ {\bm{W}}^{y}_{22}={\bm{w}}_{v_{21}}{\bm{w}}_{y_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{y_{21}}^{\rm T},\\ \breve{\bm{W}}^{y}_{11}={\bm{w}}_{v_{12}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{y_{12}}^{\rm T},\\ \breve{\bm{W}}^{y}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{y_{11}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{y_{12}}^{\rm T}.\end{cases}$ (18)
Similarly, the velocity CRLB matrix can be rewritten as
$\small{\bm{C}}_{\rm V}=\frac{1}{|{\bm{H}}|}\begin{bmatrix}h_{22}&-h_{12}\\ -h_{21}&h_{11}\end{bmatrix},$ (19)
where $|{\bm{H}}|=h_{11}h_{22}-h_{12}h_{21}$. The relevant elements $h_{ij}$ are given respectively by
$\small h_{11}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{11}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{21}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$ (20)
$h_{12}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{12}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{11}^{p}){\bm{\rho}}\big{]},$
$h_{22}=\frac{1}{|{\bm{P}}|}\big{[}{\bm{\rho}}^{\rm T}({\bm{w}}_{y_{22}}{\bm{\rho}}^{\rm T}{\bm{W}}_{p}\!-\!{\bm{w}}_{v_{12}}{\bm{\rho}}^{\rm T}{\bm{W}}_{22}^{p}\!-\!{\bm{w}}_{v_{22}}{\bm{\rho}}^{\rm T}\breve{{\bm{W}}}_{21}^{p}){\bm{\rho}}\big{]},$
where $|{\bm{P}}|={\bm{\rho}}^{\rm T}{\bm{W}}_{p}{\bm{\rho}}$, and
$\small\begin{cases}{\bm{W}}_{p}={\bm{w}}_{p_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{p_{12}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{12}={\bm{w}}_{v_{11}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{21}}{\bm{w}}_{p_{21}}^{\rm T},\\ {\bm{W}}^{p}_{22}={\bm{w}}_{v_{12}}{\bm{w}}_{p_{22}}^{\rm T}-{\bm{w}}_{v_{22}}{\bm{w}}_{p_{21}}^{\rm T},\\ \breve{\bm{W}}^{p}_{11}={\bm{w}}_{v_{21}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{11}}{\bm{w}}_{p_{12}}^{\rm T},\\ \breve{\bm{W}}^{p}_{21}={\bm{w}}_{v_{22}}{\bm{w}}_{p_{11}}^{\rm T}-{\bm{w}}_{v_{12}}{\bm{w}}_{p_{12}}^{\rm T}.\end{cases}$ (21)
Note that the newly derived CRLB expressions in (16) and (19) represent explicit functions of the PA vector, thereby serving as effective and tractable sensing performance metrics.
III-B Problem Formulation
This section formulates the PA problem by maximizing the overall performance of CF MIMO-ISAC systems. Depending on specific application requirements, the optimization objective can be either sensing or communication performance, while the requirement of the other functionality is incorporated as a constraint in the optimization model.
Let ${\bm{\rho}}_{\min}=[\rho_{\min,1},\rho_{\min,2},\cdots,\rho_{\min,N}]^{\rm T}$ denote the minimum PA coefficient to achieve the sufficiently large SNR for tight CRLBs [34, 36], and ${\bm{\rho}}_{\max}=[\rho_{\max,1},\rho_{\max,2},\cdots,\rho_{\max,N}]^{\rm T}$ denote the maximum PA coefficient due to the power budgets of individual transmitters [38, 36]. In this paper, we consider the problem of maximizing the communication SINR, subject to constraints on the sensing performance, i.e.,
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ -{\bm{\rho}}^{\rm T}{\bm{g}},$ (22a)
${\rm s.t.}$ ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}=1,$ (22b)
${\bm{\rho}}_{\min}\leq{\bm{\rho}}\leq{\bm{\rho}}_{\max},$ (22c)
${\rm tr}\{{\bm{C}}_{\rm L}\}\leq\delta_{l}^{2},$ (22d)
${\rm tr}\{{\bm{C}}_{\rm V}\}\leq\delta_{v}^{2},$ (22e)
where $\delta_{l}^{2}$ and $\delta_{v}^{2}$ denote the corresponding location and velocity CRLB thresholds.
IV Power Allocation Algorithm for CF MIMO-ISAC
In this section, we investigate the optimal PA strategy for maximizing the ISAC performance, as formulated in (22a)-(22e). Unfortunately, the strong nonlinearity and non-convexity of the CRLB-based constraints in (22d) and (22e) render the problem analytically intractable. To overcome this challenge, we adopt a CG iterative algorithm, enhanced with a penalty function and projection operator, to provide an efficient numerical solution.
To derive a numerical solution, we start from converting the original constrained optimization problem into an unconstrained one. To this end, we first introduce a penalty (function) for violating the nonlinear inequality constraints in (22c)-(22e), defined as [39]
$\alpha({\bm{\rho}})=\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})+\sum\nolimits_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$
where the involved penalty sub-functions are given respectively by
$\small\begin{cases}\tilde{f}_{l}({\bm{\rho}})=(\max\{0,{f}_{l}({\bm{\rho}})\})^{q},\\ \tilde{f}_{v}({\bm{\rho}})=(\max\{0,{f}_{v}({\bm{\rho}})\})^{q},\\ f_{\min,n}({\bm{\rho}})=(\max\{0,\rho_{\min,n}-\rho_{n}\})^{q},\\ f_{\max,n}({\bm{\rho}})=(\max\{0,\rho_{n}-\rho_{\max,n}\})^{q},\end{cases}$ (23)
with ${f}_{l}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm L}\}-\delta_{l}^{2}$ and ${f}_{v}({\bm{\rho}})\triangleq{\rm tr}\{{\bm{C}}_{\rm V}\}-\delta_{v}^{2}$. $q$ is the power of the penalty function. We consider $q=2$ to guarantee their continuous differentiability. Then, the objective function for the penalty problem can be written as
$\small L({\bm{\rho}})=-{\bm{\rho}}^{\rm T}{\bm{g}}+\mu\alpha({\bm{\rho}}),$ (24)
where $\mu$ denotes the penalty factor.
Additionally, to enforce this equality constraint (22b) and reduce both computational complexity and the nonlinearity of the optimization problem, we adopt a projection operator instead of incorporating another penalty sub-function, i.e.,
$\small{\bm{\rho}}({\bm{\varpi}})={\bm{\Theta}}{\bm{\varpi}}+{\bm{\rho}}_{0},$ (25)
where ${\bm{\varpi}}\in\mathbb{R}^{N\times 1}$ is an arbitrary vector and ${\bm{\Theta}}\triangleq{\bm{I}}_{N\times N}-\frac{{\bm{1}}_{N}{\bm{1}}_{N}^{\rm T}}{N}$ is a projection matrix onto the zero space of ${\bm{1}}_{N}$. We have ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Thus, the equality constraint ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}({\bm{\varpi}})=1$ in (22b) always holds by initializing ${\bm{\rho}}_{0}$ such that ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{0}=1$. Consequently, ${\bm{\rho}}({\bm{\varpi}})$ is the sum of a constant PA vector ${\bm{\rho}}_{0}$ and a projection of ${\bm{\varpi}}$ onto the zero space of ${\bm{1}}_{N}$.
By substituting (25) into (24), we can convert the original constrained optimization to an unconstrained optimization problem as follows:
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ \mathcal{L}({\bm{\varpi}})=-{\bm{\rho}}^{\rm T}({\bm{\varpi}}){\bm{g}}+\mu\alpha({\bm{\rho}}({\bm{\varpi}})).$ (26)
Then, unconstrained optimization methods, e.g., CG and SD methods, can be employed to search for the optimal solution.
IV-A Modified Conjugate Gradient Algorithm
To numerically solve (26), we propose a projection-based MCG algorithm. The update at the $i$th iteration is given by
$\small{\bm{\varpi}}_{i+1}={\bm{\varpi}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (27)
where ${\bm{d}}_{i}$ is the search direction of $\mathcal{L}({\bm{\varpi}}_{i})$ and $\upsilon_{i}$ is the step size. The search direction using the MCG is updated by
${\bm{d}}_{i+1}$ $=-\nabla\mathcal{L}({\bm{\varpi}}_{i+1})+\varsigma_{i}{\bm{d}}_{i},$ (28)
where $\varsigma_{i}$ denotes the deflection factor of a CG algorithm. Taking the derivative $\mathcal{L}({\bm{\varpi}})$ regarding ${\bm{\varpi}}$ yields
(ϖ)=bold-italic-$\small\nabla\mathcal{L}({\bm{\varpi}})=$ ${\bm{\Theta}}\nabla L({\bm{\rho}})={\bm{\Theta}}\big{[}{\bm{g}}+\nabla\alpha({\bm{\rho}})\big{]},$ (29)
where
$\small\nabla\alpha({\bm{\rho}})=$ $\mu\big{(}\nabla\tilde{f}_{l}({\bm{\rho}})+\nabla\tilde{f}_{v}({\bm{\rho}})$
$+\sum\nolimits_{n=1}^{N}(\nabla f_{\min,n}({\bm{\rho}})+\nabla f_{\max,n}({\bm{\rho}}))\big{)},$ (30)
and
$\small\begin{cases}\nabla\tilde{f}_{l}({\bm{\rho}})=2\max\{f_{l}({\bm{\rho}}),0\}\nabla f_{l}({\bm{\rho}})\\ \nabla\tilde{f}_{v}({\bm{\rho}})=2\max\{f_{v}({\bm{\rho}}),0\}\nabla f_{v}({\bm{\rho}})\\ \nabla f_{\min,n}({\bm{\rho}})=2\min\{\rho_{n}-\rho_{\min,n},0\}{\bm{e}}_{n}\\ \nabla f_{\max,n}({\bm{\rho}})=2\max\{\rho_{n}-\rho_{\max,n},0\}{\bm{e}}_{n}.\end{cases}$ (31)
Note that ${\bm{e}}_{n}\in\mathbb{R}^{N\times 1}$ is an index vector with its element at index $n$ being $1$ and the others being zeros.
Given the initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$, and using the idempotent property of the projection matrix, i.e., ${\bm{\Theta}}^{2}={\bm{\Theta}}$, we have ${\bm{\Theta}}{\bm{d}}_{i}={\bm{d}}_{i}$, $\forall{i>0}$. Thus, taking the projection of (27) leads to
$\small{\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i},$ (32)
where the search direction is updated by
$\small{\bm{d}}_{i+1}$ $=-{\bm{\Theta}}\nabla\mathcal{L}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}.$ (33)
We have ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{i}=0$ for $\forall{i>1}$ due to ${\bm{1}}_{N}^{\rm T}{\bm{d}}_{1}=0$ and ${\bm{1}}_{N}^{\rm T}{\bm{\Theta}}={\bm{0}}_{N}^{\rm T}$. Therefore, the equality constraint in (22b), that is, ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{i}=1$, always holds at $i$th iteration for $\forall{i>1}$ if and only if the initial PA vector ${\bm{\rho}}_{1}$ satisfies ${\bm{1}}_{N}^{\rm T}{\bm{\rho}}_{1}=1$.
We now investigate the selection of the deflection factor $\varsigma_{i}$ and step size $\upsilon_{i}$ at each iteration, which is significant in guaranteeing the convergence of the MCG algorithm. One necessary condition for the convergence is the non-increasing objective function at each iteration, i.e.,
$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq 0.$ (34)
To this end, the search direction ${\bm{d}}_{i}$ should be the descent direction and the step size $\upsilon_{i}$ should be correspondingly upper bounded by a threshold at each iteration.
IV-A1 Descent direction
To ensure a descent direction ${\bm{d}}_{i+1}$ at the $(i+1)$th iteration, it requires
$\small{\bm{d}}_{i+1}^{\rm T}\nabla L({\bm{\rho}}_{i+1})\leq 0,$ (35)
where the equality holds if and only if $\nabla L({\bm{\rho}}_{i+1})={\bm{0}}$. The initial direction ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$ is the descent direction of $L({\bm{\rho}}_{1})$ with $-\nabla L({\bm{\rho}}_{1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\leq 0$ since the projection operator ${\bm{\Theta}}$ is a symmetric positive definite matrix. Substituting (33) into (35), we have
$\small\begin{cases}\varsigma_{i}\leq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}>0\\ \varsigma_{i}\geq\varsigma_{i}^{\rm b},\ \varsigma_{i}^{\rm b}<0\end{cases}$ (36)
where $\varsigma_{i}^{\rm b}$ is the bound for $\varsigma_{i}$ with its expression given by
$\small\varsigma_{i}^{\rm b}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})}{{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})}.$ (37)
The numerator satisfies $\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})>0$ for any non-zero gradient $\nabla L({\bm{\rho}}_{i+1})$ due to the positive definite ${\bm{\Theta}}$.
When employing inexact line searches, the deflection selection proposed by Hestenes and Stiefel [40], referred to as HS method, for $\varsigma_{i}$ is preferable. Accordingly, we adopt the modified HS deflection, given by
$\small\varsigma_{i}^{\rm HS}=\dfrac{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}}{{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}},$ (38)
which is obtained from ${\bm{d}}_{i+1}^{\rm T}{\bm{q}}_{i}=0$ with ${\bm{q}}_{i}\triangleq{\bm{\Theta}}[\nabla L({\bm{\rho}}_{i+1})-\nabla L({\bm{\rho}}_{i})]$. In fact, when using the MCG algorithm, it is desirable to maintain orthogonality between the conjugate gradients of adjacent iterations, i.e.,
$|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})|,$ (39)
$|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i})|\ll|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|.$ (40)
(39) and (40) ensure the positivity of ${\bm{d}}_{i}^{\rm T}{\bm{q}}i$ and $\nabla L({\bm{\rho}}{i+1})^{\rm T}{\bm{q}}_{i}$, respectively. This allows for a positive conjugate gradient scaling factor $\varsigma_{i}^{\rm HS}$. Furthermore, (39) and (40) also imply that $|{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i+1})|\ll|{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i}|$ and $|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i}|\approx|\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})|$, respectively. As a result, $\varsigma_{i}^{\rm HS}$ is expected to be smaller than $|\varsigma_{i}^{\rm b}|$, thereby satisfying the deflection constraint defined in (36).
Rather than strictly relying on the potentially tight and unquantifiable conjugacy conditions in (39) and (40), we instead enforce the positivity of both the numerator and denominator in (38), along with the inequality $\varsigma_{i}^{\rm HS}<|\varsigma_{i}^{\rm b}|$. When these conditions hold, we set $\varsigma_{i}=\varsigma_{i}^{\rm HS}$. Otherwise, we reset the MCG iteration by letting $\varsigma_{i}=0$. To sum up, the deflection factor is updated by
${\varsigma_{i}=}\begin{cases}\varsigma_{i}^{\rm HS},\ \min\{\nabla L({\bm{\rho}}_{i+1})^{\rm T}{\bm{q}}_{i},{\bm{d}}_{i}^{\rm T}{\bm{q}}_{i},|\varsigma_{i}^{\rm b}|-\varsigma_{i}^{\rm HS}\}>0,\\ 0,\ {\rm else}.\end{cases}$ (41)
Note that, after every consecutive $N$ MCG steps with $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, we reset $\varsigma_{i}=0$ to initiate a new cycle of MCG iterations. Such a reboot strategy can help mitigate the accumulation of numerical errors, restoring the orthogonality and conjugacy of direction vectors, and avoiding local minima in nonlinear and non-convex optimization.
IV-A2 Step size optimization
To optimize the step size selection, we define $f_{i}(\upsilon_{i})\triangleq L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})$ with $f_{i}(0)=L({\bm{\rho}}_{i})$. The second-order Taylor approximation of $f_{i}(\upsilon_{i})$ at $\upsilon_{i}=0$ is given by
$\small f_{i}(\upsilon_{i})$ $\thickapprox f_{i}(0)+\upsilon_{i}\nabla f_{i}(0)+\upsilon_{i}^{2}\nabla^{2}f_{i}(0)/2$
$=f_{i}(0)+\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i},$ (42)
where $\nabla^{2}L({\bm{\rho}}_{i})$ is the Hessian matrix of $L({\bm{\rho}})$ at ${\bm{\rho}}_{i}$. According to (35), we have ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})\leq 0$. If $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is locally bounded, there exists a non-negative step size bound $\upsilon_{b}$ for given $\epsilon_{L}\in[0,1)$ such that
$\small f_{i}(\upsilon_{i})-f_{i}(0)\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}),\quad\text{for}\quad\upsilon_{i}\leq\upsilon_{b}.$ (43)
(43) always holds when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq 0$. In contrast, we have $\upsilon_{b}=\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}$ when ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}>0$.
Due to the inherent high computational complexity of the Hessian matrix $\nabla^{2}L({\bm{\rho}}_{i})$ in calaulating $\upsilon_{b}$, we can apply an inexact line search (ILS) for the step size to enable
$\small L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}).$ (44)
According to (35), the term on the right-hand side of (44) is always non-positive. Thus, (44) can enable a descent function in (34). The ILS controlling factor $\epsilon_{L}$ is used to regulate the decrement of the function $L({\bm{\rho}})$ by moving ${\bm{\rho}}_{i}$ towards the direction ${\bm{d}}_{i}$ weighted by the step size $\upsilon_{i}$. When (44) is unmet for given $\upsilon_{i}$, the step size is scaled by
$\small\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i},$ (45)
where $\sigma_{\upsilon}\in(0,1)$ denotes the scaling factor and an initial step size is denoted as $\upsilon_{i}=\upsilon_{0,i}$.
Additionally, the penalty function does not inherently ensure that the individual power budget constraints at each transmitter are respected during the optimization process. As a result, negative PA values may arise in intermediate iterations. These infeasible PA values can lead to misleading CRLB evaluations, for instance, producing negative CRLBs instead of infinite ones, thereby falsely satisfying the CRLB constraints. To avoid unnecessary iterations within the infeasible region and to ensure physically meaningful results, it is crucial to enforce non-negative PAs at each iteration step. To this end, the step size $\upsilon_{i}$ needs to satisfy
$\small 0<\upsilon_{i}\leq\upsilon_{i}^{\rm b},$ (46)
where $\upsilon_{i}^{\rm b}$ denotes an upper bound on the step size, given by
$\upsilon_{i}^{\rm b}=\min_{d_{n,i}<0}\{\rho_{n,i}/|{d}_{n,i}|\},$ (47)
with $\rho_{n,i}$ denoting the $n$th element of ${\bm{\rho}}_{i}$ and similar for ${d}_{n,i}$. Thus, the initial step size for ILS can be selected as
$\small\upsilon_{0,i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$ (48)
with $\epsilon_{\upsilon}\in(0,1]$. Note that there must exist at least one component ${d}_{n,i}<0$ in each iteration $i$, unless ${\bm{d}}_{i}={\bm{0}}$, due to the constraint $\sum_{n=1}^{N}{d}_{n,i}=0$. The ILS method characterized in (44)-(48) can be seen as the extension and application of the Armijo’s Rule [41] in our work.
IV-A3 The algorithm implementation
A well-defined termination condition is essential for the proposed iterative algorithm. A natural choice, inspired by the ILS condition in (44), is to monitor the absolute change in the objective function between consecutive iterations, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|\leq\epsilon$, where $\epsilon$ is a predefined threshold with small value. To avoid numerical instability arising from the magnitude of the objective function itself, we apply a normalized termination condition, i.e., $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$. This condition is employed in both the MCG-based and MSD-based algorithms.
We now present the complete PP-MCG-ILS. Generally, a large penalty factor may lead to numerical issues, such as an ill-conditioned Hessian matrix, which can degrade the convergence behavior of the MCG algorithm. To mitigate this, we adopt a sequential penalty strategy that gradually increases the penalty factor across successive iteration loops. Specifically, for a given penalty factor $\mu_{k}$, the MCG algorithm is executed until its termination condition is satisfied. Then, the penalty factor is updated as $\mu_{k+1}=\varphi\mu_{k}$ where $\varphi>1$ is a scaling constant, for the next loop of penalty-based iteration. This iterative refinement continues until the convergence criterion $\mu_{k}\alpha(\bm{\rho})\leq\epsilon_{\mu}$ is satisfied, where $\epsilon_{\mu}$ is a small, predefined threshold [39]. This sequential penalty strategy ensures both numerical stability and progressive enforcement of the constraint conditions. The detailed procedure of the proposed PP-MCG-ILS algorithm is summarized in Algorithm 1.
Algorithm 1 The PP-MCG-ILS algorithm
Input: The positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constant $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial penalty factor $\mu_{1}$, scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
2:  Step size search: set $\upsilon_{i}=\epsilon_{\upsilon}\upsilon_{i}^{\rm b}$, while $L({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
4:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 7; Otherwise, go to next step.
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
6:  Direction update: ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
IV-B Modified Steepest Descent Algorithm
In fact, alternative unconstrained optimization methods can be applied to solve the proposed projection-based penalty problem in (26). In this paper, we also develop a MSD algorithm, adapted from the proposed MCG framework. The main difference between the MSD and MCG algorithms lies in how the search direction is determined. In the MSD approach, the search direction at each iteration is set to the negative gradient of the objective function, i.e.,
$\small{\bm{d}}_{i}$ $=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i}),$ (49)
which can be obtained from MCG by setting the deflection factor to zero. We herein omit the detailed implementation steps of PP-MSD-ILS since they can be obtained from the PP-MCG-ILS in Algorithm 1 by simply removing step 5 and performing the direction update in step 6 using (49).
IV-C Benchmarking Schemes
To facilitate performance comparison, we introduce two benchmark algorithms based on the PP-MCG-ILS and PP-MSD-ILS frameworks, called penalty function and projection-driven normalized SD (PP-NSD) and penalty function and projection-driven normalized CG (PP-NCG), respectively. Here, we employ the standard SD and CG methods proposed in [42], where fixed step sizes are adopted instead of ILS. To mitigate potential instability arising from overly large or small gradient magnitudes under fixed-step updates, we normalize the update directions using their $l_{2}$ norm. Specifically, for the PP-NSD algorithm, the search direction is defined as an $\epsilon_{d}$-regularized normalized gradient, given by
${\bm{d}}_{i}$ $=-{\bm{g}}_{i}/(\|{\bm{g}}_{i}\|_{2}+\epsilon_{d}),$ (50)
where ${\bm{g}}_{i}={\bm{\Theta}}\nabla L({\bm{\rho}}_{i})$, and $\epsilon_{d}$ is a small positive constant introduced to prevent division by zero. Similarly, for the PP-NCG algorithm, the search direction is given by an $\epsilon_{d}$-regularized normalized conjugate gradient, i.e.,
$\small{\bm{d}}_{i+1}$ $=-({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})/(\|({\bm{g}}_{i+1}+\varsigma_{i}{\bm{d}}_{i})\|_{2}+\epsilon_{d}),$ (51)
where the deflection factor $\varsigma_{i}$ is computed by (41).
The step size in both benchmark algorithms is given by
$\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\},$ (52)
with $\upsilon_{i}^{\rm b}$ given in (47) and $\upsilon_{0}$ being a predetermined step size. The maximum number of iterations is set to $I$. Due to the non-monotonic penalty function, the minimum may not occur at the final iteration but within a few steps before it. To address this, we introduce an alternative penalty stopping criterion given by $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$ where $S$ represents the window size for searching the minimum penalty value. The detailed implementation steps of the PP-NSD algorithm are omitted here, as it can be easily constructed from the PP-NCG described in Algorithm 2 by removing step 4 and performing the direction update in step 5 using (50).
Algorithm 2 The PP-NCG algorithm
Input: Positive scalar termination factor $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, a large integer $I$ and a small integer $S$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$, $\epsilon_{d}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, projection operator ${\bm{\Theta}}$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{1})\|_{2}+\epsilon_{d})$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$.
2:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
3:  Convergence condition: If $|L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i})|/|L({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$ or $i+1=I$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, go to step 6; Otherwise, go to next step.
4:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$;
5:  Direction update: ${\bm{d}}_{i+1}$ through (51); let $i=i+1$, and go back to step 1.
6:  Penalty update: If $\mu_{j}\min\limits_{i+1-S\leq q\leq i+1}\{\alpha({\bm{\rho}}_{q})\}<\epsilon_{\mu}$, stop; Otherwise, let $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})/(\|{\bm{\Theta}}\nabla L({\bm{\rho}}_{i+1})\|_{2}+\epsilon_{d})$; let $j=j+1$ and $i=i+1$, and go to step 1.
V Extension to Pure Sensing
In this section, we consider a pure sensing scenario with joint location and velocity estimation. The P-NCG-ILS algorithm is proposed for PA in minimizing the total power under power budget constraints and CRLB constraints.
The constrained power minimization problem is formulated as
$\small\arg\min\limits_{{\bm{\rho}}}$ $\ {\bm{\rho}}^{\rm T}{\bm{1}},$ (53a)
${\rm s.t.}$ ${\bm{\rho}}^{\rm T}{\bm{1}}\leq 1,$ (53b)
[…]
Analogously to the derivations of PP-MCG-ILS, we set the penalty function for violating the inequality constraints as
$\small\alpha_{\rm s}({\bm{\rho}})=$ ${f}_{\rm s}({\bm{\rho}})+\tilde{f}_{l}({\bm{\rho}})+\tilde{f}_{v}({\bm{\rho}})$
$+\sum_{n=1}^{N}\left(f_{\min,n}({\bm{\rho}})+f_{\max,n}({\bm{\rho}})\right),$ (54)
where ${f}_{\rm s}({\bm{\rho}})$ denotes the penalty sub-function for the total power budget, i.e.,
$\small{f}_{\rm s}({\bm{\rho}})=(\max\{0,{\bm{\rho}}^{\rm T}{\bm{1}}-1\})^{q}.$ (55)
Then, the unconstrained optimization problem can be formulated as
$\small\arg\min\limits_{{\bm{\rho}}}\ L_{\rm s}({\bm{\rho}})={\bm{\rho}}^{\rm T}{\bm{1}}+\mu\alpha_{\rm s}({\bm{\rho}}).$ (56)
Similarly to the PP-MCG-ILS algorithm, we also use the CG-based method to solve (56). To avoid a potential numerical instability, we consider normalized CG search direction, given by ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$, where the deflection computation method is the same as that of the PP-MCG-ILS algorithm. Additionally, ILS is used to determine the step size at each iteration. The complete P-NCG-ILS algorithm is summarized in Algorithm 3.
Algorithm 3 The P-NCG-ILS algorithm
Input: Positive scalar termination factors $\epsilon_{\mu}$ and $\epsilon_{\rm th}$, small positive constants $\sigma_{\upsilon}$, $\epsilon_{L}$ and $\epsilon_{\upsilon}$, initial step size $\upsilon_{0}$, initial penalty factor $\mu_{1}$ and its scaling factor $\varphi>1$; initial PA parameter ${\bm{\rho}}_{1}$ satisfying ${\bm{1}}^{\rm T}{\bm{\rho}}_{1}=1$, ${\bm{y}}_{1}={\bm{\rho}}_{1}$, and ${\bm{d}}_{1}=-\nabla L_{\rm s}({\bm{\rho}}_{1})/\|\nabla L_{\rm s}({\bm{\rho}}_{1})\|_{2}$; a counter $k=0$ and $i=j=1$.
Output: The PA vector ${\bm{y}}_{j}$.
1:  Step size bound: Calculate $\upsilon_{i}^{\rm b}$ according to (47);
2:  Step size: $\upsilon_{i}=\min\{\epsilon_{\upsilon}\upsilon_{i}^{\rm b},\upsilon_{0}\}$; while $L_{\rm s}({\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i})-L_{\rm s}({\bm{\rho}}_{i})>\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$, let $\upsilon_{i}=\sigma_{\upsilon}\upsilon_{i}$.
3:  PA update: ${\bm{\rho}}_{i+1}={\bm{\rho}}_{i}+\upsilon_{i}{\bm{d}}_{i}$.
4:  Convergence condition: If $|L_{\rm s}({\bm{\rho}}_{i+1})-L_{\rm s}({\bm{\rho}}_{i})|/|L_{\rm s}({\bm{\rho}}_{i+1})|\leq\epsilon_{\rm th}$, ${\bm{y}}_{j}={\bm{\rho}}_{i+1}$, and go to step 5; Otherwise, go to next step.
5:  Deflection choice: According to (41), if $\varsigma_{i}=\varsigma_{i}^{\rm HS}$, $k=k+1$, or $k=0$ otherwise; when $k=N$, let $\varsigma_{i}=0$ and $k=0$.
6:  Direction update: ${\bm{d}}_{i+1}=\tilde{\bm{d}}_{i+1}/(\|\tilde{\bm{d}}_{i+1}\|_{2}+\epsilon_{d})$ with $\tilde{\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})+\varsigma_{i}{\bm{d}}_{i}$; let $i=i+1$, and go back to step 1.
7:  Penalty update: If $\mu_{j}\alpha({\bm{y}}_{j})<\epsilon_{\mu}$, stop; Otherwise, $\mu_{j+1}=\varphi\mu_{j}$, and ${\bm{d}}_{i+1}=-\nabla L_{\rm s}({\bm{\rho}}_{i+1})$; let $j=j+1$ and $i=i+1$, and go to step 1.
VI Algorithm Analysis
The proposed algorithms (e.g., PP-MCG-ILS, PP-MSD-ILS) can be extended and applied to a broad class of nonlinear and non-convex optimization problems. These problems are typically characterized by an affine equality constraint, if present, satisfied through a projection operator, along with nonlinear or non-convex constraints incorporated via penalty functions. The ILS mechanism is used to enhance convergence. For example, the proposed P-NCG-ILS algorithm for sensing-only scenario can be seen as a simple extension of the PP-MCG-ILS algorithm. In the following, we analyze the convergence behavior and computational complexity of the proposed algorithms.
VI-A Convergence Analysis
We now provide the convergence analysis of the proposed PP-MCG-ILS and PP-MSD-ILS algorithms. Before delving to the details, we define the increment of the parameter update and the objective function respectively as
$\small\Delta{\bm{\rho}}_{i+1}$ $\triangleq{\bm{\rho}}_{i+1}-{\bm{\rho}}_{i}=v_{i}{\bm{d}}_{i},$ (57)
and
$\Delta L({\bm{\rho}}_{i+1})\triangleq L({\bm{\rho}}_{i+1})-L({\bm{\rho}}_{i}).$ (58)
Lemma 1. The sufficient and necessary condition for $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ is that, there exists an integer $I$ such that, ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$.
Proof. The proof is given in Appendix B. $\blacksquare$
Definition. A function $L({\bm{\rho}})$ is Lipschitz gradient continuous over a set $\Omega$, if there exists a constant $C>0$, $\|\nabla L({\bm{\rho}}_{i})-\nabla L({\bm{\rho}}_{j})\|\leq C\|{\bm{\rho}}_{i}-{\bm{\rho}}_{j}\|$ for any ${\bm{\rho}}_{i},{\bm{\rho}}_{j}\in\Omega$ [43].
Theorem 1. Given a nonempty convex set $\Omega$ with the objective function $L({\bm{\rho}}_{i})$ bounded over its interior points, if the objective function is Lipschitz gradient continuous over $\Omega$, the ILS-based iterative algorithm will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$. In contrast, if it is Lipschitz gradient non-continuous, sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ may converge to a Lipschitz gradient noncontinuous point with the norm of the Hessian matrix approaching infinity.
Proof. The proof is given in Appendix C. $\blacksquare$
Theorem 1 provides the convergence analysis for the proposed iterative algorithms.
We now prove the Lipschitz gradient continuity of the objectivity functions in (26) and (56).
Firstly, all CRLB-involved matrices, e.g., ${\bm{P}}(\rho)$, ${\bm{Y}}(\rho)$, and their inverses are smooth on ${\bm{\rho}}$ over the feasible domain $\Omega\subset\mathbb{R}^{N}$. Moreover, ${\bm{C}}_{L}({\bm{\rho}})$ is uniformly positive definite. It follows that ${\bm{C}}_{L}({\bm{\rho}})$ is continuously differentiable, and so is ${\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\}$. By the chain rule, we have
$\nabla\tilde{f}_{l}({\bm{\rho}})=\begin{cases}2f_{l}({\bm{\rho}})\nabla f_{l}({\bm{\rho}}),&\text{if }f_{l}({\bm{\rho}})>0\\ 0,&\text{otherwise}.\end{cases}$ (59)
where $\nabla f_{l}({\bm{\rho}})=-{\rm tr}\{{\bm{C}}_{L}({\bm{\rho}})\nabla({\bm{P}}-{\bm{V}}{\bm{Y}}^{-1}{\bm{V}}^{\rm T}){\bm{C}}_{L}({\bm{\rho}})\}$. Since the matrix inverse and trace are smooth on the set of positive definite matrices, $\nabla f_{l}({\bm{\rho}})$ is continuously differentiable. Moreover, all matrices involved (including ${\bm{C}}_{L}({\bm{\rho}})$) have bounded spectral norm, so the operator norm of the Hessian of $f_{l}({\bm{\rho}})$ is also bounded. Consequently, $\nabla\tilde{f}_{l}({\bm{\rho}})$ is Lipschitz continuous. A similar argument applies to $\nabla\tilde{f}_{v}({\bm{\rho}})$. On the other hand, $f_{\min,n}({\bm{\rho}})$, $f_{\max,n}({\bm{\rho}})$ and ${f}_{\rm s}({\bm{\rho}})$ are piecewise quadratic, and their gradients are Lipschitz continuous. Hence, $\nabla_{\bm{\rho}}\alpha({\bm{\rho}})$ and $\nabla_{\bm{\rho}}\alpha_{\rm s}({\bm{\rho}})$ are Lipschitz continuous over $\Omega$, and the same applies to $\nabla_{\bm{\rho}}L({\bm{\rho}})$ and $\nabla_{\bm{\rho}}L_{\rm s}({\bm{\rho}})$.
Additionally, since ${\bm{\Theta}}$ is a constant matrix with $\|{\bm{\Theta}}\|_{2}=1$, and $\nabla_{\bm{\rho}}f({\bm{\rho}})$ is Lipschitz continuous, the composition $\nabla\mathcal{L}({\bm{\varpi}})={\bm{\Theta}}\nabla L({\bm{\rho}})$ is also Lipschitz continuous. In other words, there exists a constant $C>0$ such that
$\small\|\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{1})-\nabla_{{\bm{\varpi}}}\mathcal{L}({\bm{\varpi}}_{2})\|_{2}\leq C\|{\bm{\varpi}}_{1}-{\bm{\varpi}}_{2}\|_{2},\forall{\bm{\varpi}}_{1},{\bm{\varpi}}_{2}\in\Omega_{\bm{\varpi}}$
Therefore, $\mathcal{L}({\bm{\varpi}})$ is Lipschitz gradient continuous on $\Omega_{\bm{\varpi}}$. $\blacksquare$
VI-B Complexity Analysis
We now qualitatively analyze and compare the computational complexity of the proposed algorithms and their benchamrks.
We observe that the increased computations of the PP-MCG-ILS algorithm, compared to the PP-MSD-ILS algorithm, primarily stems from the computation of $\varsigma_{i}^{\rm b}$ and $\varsigma_{i}^{\rm HS}$ at each iteration, which involves approximately $4N$ multiplications and 2 divisions. In both algorithms, the dominant computational burden lies in evaluating the objective function during the ILS and computing the gradients for direction updates, which are inherently tied to the CRLB expressions. Therefore, a more meaningful comparison of the overall computational cost focuses on the number of ILS steps performed per iteration. This number is influenced by the step size threshold, given by the right-hand side of (C), the initial step size $\upsilon_{0,i}$ in (48), and the scaling factor. At a given iterate ${\bm{\rho}}_{i}$, a larger search direction norm $\|{\bm{d}}_{i}\|$ generally leads to a smaller $\upsilon_{0,i}$ and a reduced step size threshold. As a consequence, the number of ILS steps per iteration is expected to remain comparable between the PP-MCG-ILS and PP-MSD-ILS algorithms. Accordingly, the primary metrics for evaluating the relative performance of these two algorithms are their convergence rate and the quality of their steady-state solutions.
We now qualitatively compare the computational complexity between algorithms employing ILS and those without ILS. As previously discussed, the number of ILS steps significantly impacts the number of evaluations of the objective function. Thus, the computational complexity for step size selection of the benchmark algorithms can be considerably lower at each iteration compared to the PP-MCG-ILS and PP-MSD-ILS algorithms. However, all four algorithms have similar complexity in direction update due to the inherently intensive gradient calculations involved. Moreover, the overall computational complexity heavily depends on the convergence rates, or equivalently, on the total number of iterations required for convergence. As will be demonstrated in the subsequent simulation results, the proposed PP-MCG-ILS algorithm exhibits a notably higher convergence rate compared to its counterparts, with or without ILS. This superior convergence performance significantly reduces the overall computational effort, particularly in the sense that it benefits demanding gradient computations.
VII Simulation Results
In this section, we conduct numerical simulations to evaluate the proposed PP-MCG-ILS, PP-MSD-ILS and P-NCG-ILS algorithms for PA optimization. All APs are assumed to be stationary. We employ the orthogonal chirp division multiplexing (OCDM) [44] waveform with a single Gaussian pulse shaping. The carrier frequency is $3$GHz. The lowpass equivalent is formulated by
$s_{n}(t)=(2/T^{2})^{1/4}e^{-\pi t^{2}/T^{2}}e^{j\pi M/T^{2}(t-(n-1)T/M)^{2}},$ (60)
where $M$ is the number of OCDM chirps, and $T$ is proportional to the effective pulse time width.
VII-A CRLB Validation
We first verify the CRLB in comparison with the MLE results. To maintain consistency with existing works [28, 38, 29], we use the signal energy-to-noise ratio (SENR), defined as $\tilde{\delta}_{\rm cn}\triangleq P/\sigma_{\rm cn}^{2}$, instead of the SCNR $\delta_{\rm cn}$ in evaluating the CRLBs. Due to the high complexity of four-dimensional MLE, we consider a simple $4\times 3$ nonsymmetric multi-static radar configuration given in Fig. 2. Without loss of generality, we consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=1$ kHz.
Fig. 3 compares the mean squared error (MSE) of the MLE with the derived CRLB. It is observed that the derived CRLBs closely match the MSEs for SENR values above $-20$ dB. The small discrepancies in the low-SENR regime are attributed to the suboptimal behavior of the likelihood function under noise-dominated conditions. Overall, these results validate the effectiveness and accuracy of the proposed CRLBs.
VII-B Power Allocation for ISAC
The performance of the proposed PA algorithms is evaluated by comparison with the PP-NCG and PP-NSD benchmark algorithms. To ensure communication quality, a larger number of APs are designated for transmitting ISAC signals, while a smaller subset is used for receiving target-reflected signals. Without loss of generality, we consider a CF MIMO-ISAC system comprising 10 transmitters and 2 receivers, as illustrated in Fig. 2.
To guarantee the CRLB effectiveness, we assume the minimum PA coefficient vector is ${\bm{\rho}}_{\min}=0.01\times{\bm{1}}_{N}$. The PA budget is ${\bm{\rho}}_{\max}=0.3\times{\bm{1}}_{N}$. The SENR is set to be $-10$ dB. Within the coherence time for RCS, a typical modulus-squared RCS matrix is given by $\begin{bmatrix}0.37&0.70&1.38&0.65&2.40&0.18&0.25&0.82&0.14&0.35\\ 0.40\!&\!2.05&1.14&0.42&0.03&0.11&2.02&1.65&1.24&3.65\end{bmatrix}^{\rm T}$. Similarly, the channel gains between the user and the transmitters are assumed to be ${\bm{g}}=[2.11,12.57,5.63,0.75,0.61,1.75,0.20,2.34,14.79,9.68]^{\rm T}$. We consider the OCDM waveform with $M=16$, $T=10^{-2}$ s and a sampling rate of $f_{s}=10$ kHz. Under the given ISAC system setup and waveform configuration, the CRLB thresholds for position and velocity are set to $[250,0.13]$.
Figure 2: The AP location distribution of the ISAC system: (a) $4\times 3$ radar setting; (b) CF MIMO-ISAC setting.
Figure 3: The MSE versus the CRLB under different SENRs.
The PA vector is initialized as the uniform PA vector ${\bm{\rho}_{1}}={\bm{1}}_{N}/N$, unless noted otherwise. The initial penalty factor is set to be $\mu_{1}=10^{4}$. We hereafter consider the ILS controlling factor $\epsilon_{L}=1e-3$, the initial step size controlling factor $\epsilon_{\upsilon}=0.9$, the penalty factor scaling coefficient $\phi=10$, and the penalty threshold $\epsilon_{\mu}=10^{-3}$, unless noted otherwise. The other simulation conditions are given in Table I.
We begin by evaluating the convergence and steady-state performance of the proposed algorithms. The SINR results corresponding to the iterative PA vectors for all four algorithms are depicted in Fig. 4. As observed, the PP-MCG-ILS algorithm exhibits the fastest convergence, followed by the PP-MSD-ILS. In contrast, the benchmark algorithms converge significantly more slowly. Furthermore, the benchmarks without ILS display noticeable numerical fluctuations, particularly under large fixed step sizes, suggesting potential instability and the risk of divergence. To ensure convergence reliability, we refrain from further increasing the step sizes in these simulations.
[…]
TABLE I: Simulation parameters
  Algorithms PP-MCG-ILS PP-MSD-ILS PP-NCG PP-NSD
  Fig. 4 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.5\}$ same as PP-MCG-ILS $\upsilon\in\{4e-5,2e-5\}$ same as PP-NCG
Figs. 5 and 6 same as Fig. 4 $\upsilon=2e-5$
Figs. 7 and 8 $\epsilon_{\rm th}=1e-12$, $\sigma_{\upsilon}=0.7$ same as PP-MCG-ILS $\upsilon=2e-5$ same as PP-NCG
Figs. 9 and 10 $\{\epsilon_{\rm th},\sigma_{\upsilon}\}=\{1e-11,0.7\}$
Figure 4: SINR convergence behavior.
Figure 5: PA coefficient: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
Figure 6: CRLB values: (a) PP-MCG-ILS; (b) PP-NCG with $\upsilon=2\times 10^{-5}$.
The sequential penalty-based algorithms above leverage the optimal solution from the previous iteration loop as the initial point for the current loop under an updated penalty factor. To further explore the impact of initialization, we consider fixing the penalty factor and comparing the ISAC performance under different PA initialization strategies. In addition to the commonly used uniform PA initialization, we introduce a heuristic PA scheme, where the initial allocation is proportional to the communication channel gain, given by ${\bm{\rho}}_{1}={\bm{g}}/\sum_{n}|g_{n}|^{2}$. The results are illustrated in Figs. 7 and 8. When comparing to sequential penalty-based counterparts in Fig. 4, the PP-MCG-ILS and PP-NCG algorithms initialized with uniform PA exhibit similar convergence behavior. However, the PP-MSD-ILS and PP-NSD algorithms show degraded convergence, highlighting the superior convergence properties of CG-based algorithms over SD-based methods under identical conditions. When employing the heuristic PA initialization, all algorithms demonstrate significantly better convergence rates compared to the uniform PA case. Moreover, both PP-NCG and PP-NSD achieve noticeable performance gains relative to their sequential penalty-based counterparts, particularly PP-NCG. In contrast, the PP-MCG-ILS algorithm shows robust and superior convergence behavior, and PP-MSD-ILS exhibits a slight degradation, relative to their sequential penalty-based counterparts.
Figure 7: SINR with initial uniform PA
Figure 8: SINR with initial heuristic PA
Figure 9: Steady-state SINRs and CRLBs versus the location CRLB thresholds: (a) SINR; (b) CRLB.
Figure 10: Steady-state SINRs and CRLBs versus the velocity CRLB thresholds: (a) SINR; (b) CRLB.
To summarize, the proposed algorithms exhibit superior convergence behavior under the sequential penalty framework, consistently outperforming the benchmark methods. While the benchmark algorithms can benefit from specific conditions, such as an empirically tuned penalty factor and a well-initialized starting point, their performance is less robust and generalizable. In contrast, the PP-MCG-ILS algorithm demonstrates consistently strong convergence and steady-state accuracy across a wide range of settings, irrespective of initialization or penalty factor selection.
We now investigate the impact of the CRLB thresholds on the performance of the proposed PP-MCG-ILS algorithm. As shown in Fig. 9, relaxing the location CRLB thresholds leads to a gradual increase in the steady-state SINR, which eventually stabilizes. This trend aligns with the behavior observed in Fig. 9, where the achieved location CRLB value closely approaches its threshold and then converge. These trends reveal the trade-off between the communication SINR and the location CRLBs. Furthermore, it is noteworthy that the velocity CRLB remains largely unaffected by variations in the location CRLB thresholds. In contrast, Fig. 10 shows that the performance of the PP-MCG-ILS algorithm exhibits negligible variation with the increased velocity CRLB threshold. The observed CRLB saturation with respect to the corresponding CRLB threshold may be an inherent outcome of the introduced power budget constraints.
Figure 11: Learning results for pure sensing: (a) total power; (b) CRLBs.
VII-C Power Allocation for Pure Sensing
We now verify the effectiveness of the proposed P-MCG-ILS algorithm for total power minimization in pure sensing scenario with joint location and velocity estimation. The initial penalty factor is set to be $\mu_{1}=1$, $\sigma_{\upsilon}=0.2$ and the penalty threshold $\epsilon_{\mu}=10^{-6}$. The other simulation conditions are the same as the PP-MCG-ILS algorithm in Fig. 4.
The iterative results for total power optimization and the corresponding CRLB performance are illustrated in Fig. 11. As seen in Fig. 11, the algorithm demonstrates rapid convergence. Under the same CRLB threshold settings, the total power consumption in the pure sensing scenario is significantly lower than that in the ISAC case. This is expected, as ISAC systems must allocate sufficient power to meet both sensing and communication requirements, particularly for maximizing communication performance. Moreover, both the location and velocity CRLBs in the pure sensing case closely align with their respective thresholds. In contrast, in the ISAC scenario (refer to Fig. 6), the velocity CRLB matches its threshold loosely. This difference arises because, in pure sensing, any enhancement in CRLB performance directly benefits from increased sensing power. However, in ISAC, PA is optimized primarily for communication performance, with CRLB constraints being satisfied, potentially loosely, within acceptable margins.
VIII Conclusions
This paper presents, for the first time, a comprehensive investigation of the power allocation problem in a CF MIMO-ISAC framework, where joint location and velocity estimation were considered as the sensing task. We employed newly derived approximate CRLBs as sensing performance metrics, alongside user receive SINR as the communication metric. To address the inherent nonlinearity and non-convexity of the resulting optimization problem, we propose the PP-MCG-ILS algorithm. Additionally, the P-NCG-ILS algorithm was further proposed to handle the power minimization problem in pure sensing scenarios. We devoted into the detailed analysis of the convergence behavior and qualitatively compare the computational complexity of all proposed methods. Simulation results confirmed the accuracy of the derived CRLBs and demonstrate the efficacy of the proposed power allocation strategies in enhancing pure sensing and overall ISAC performance. Our findings offer valuable insights into resource-efficient design for next-generation CF MIMO-ISAC networks and lay the groundwork for future extensions involving more complex sensing tasks and dynamic resource management.
Appendix A Relevant parameters of the derived CRLBs
The waveform-related parameters $a_{n,k}$, $b_{n,k}$ and $d_{n,k}$, are given by
$\small[a_{n,k},\ b_{n,k},\ d_{n,k}]=\rho_{n}\big{[}\underbrace{8\pi^{2}S\bar{f_{n}^{2}}\delta_{\rm cn}}_{\bar{a}_{n,k}},\ \underbrace{4\pi S\mathfrak{T}\{\sigma_{tf}\}\delta_{\rm cn}}_{\bar{b}_{n,k}},\ \underbrace{8\pi^{2}S\bar{t_{n}^{2}}\delta_{\rm cn}}_{\bar{d}_{n,k}}\big{]}$ (61)
where $S$ denotes the number of samples. In addition, the remaining parameters, including $\beta_{n,k}$, $\zeta_{n,k}$, $\xi_{n,k}$, $\varrho_{n,k}$, $\eta_{n,k}$, and $\kappa_{n,k}$, represent geometric spreading factors, given respectively by
$[\beta_{n,k},\zeta_{n,k}]^{\rm T}=$ $\dfrac{|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}-{\bm{l}}_{n}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}-{\bm{l}}_{k}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62a)
$[\xi_{n,k},\varrho_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\left(\dfrac{{\bm{l}}_{n}-{\bm{l}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}+\dfrac{{\bm{l}}_{k}-{\bm{l}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\right),$ (62b)
$[\eta_{n,k},\kappa_{n,k}]^{\rm T}=$ $\dfrac{f_{c}|\alpha_{n,k}|}{c}\bigg{(}({\bm{l}}_{n}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{n}-{\bm{l}})}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{n}-{\bm{l}}\|_{2}}$
$+({\bm{l}}_{k}-{\bm{l}})\dfrac{{\bm{v}}^{\rm T}({\bm{l}}_{k}-{\bm{l}})}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}^{3}}-\dfrac{{\bm{v}}}{\|{\bm{l}}_{k}-{\bm{l}}\|_{2}}\bigg{)}.$ (62c)
Appendix B Proof of Lemma 1
It is obvious that $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ if and only if either ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ when $i>I$. On the other hand, in light of (42), we have
$\small\Delta L({\bm{\rho}}_{i+1})\approx\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}.$ (63)
If ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ for $i>I$, we readily have $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. Assume there exists $i>I$, where $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ with neither ${\bm{d}}_{i}\rightarrow{\bm{0}}$ nor $v_{i}\rightarrow 0$. We can solve the step size for $\Delta L({\bm{\rho}}_{i+1})=0$ as
$\small\upsilon_{i}=\dfrac{-2{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}},$ (64)
which corresponds to the step size bound in the ILS with $\epsilon_{L}=0$. However, we require $\epsilon_{L}>0$ to perform ILS effectively. Thus, ${\bm{d}}_{i}\rightarrow{\bm{0}}$ or $v_{i}\rightarrow 0$ are necessary for $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$. $\blacksquare$
Appendix C Proof of Theorem 1
In light of (42), (44) and (63), the ILS condition requires
$\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})+\upsilon_{i}^{2}/2{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}\leq\epsilon_{L}\upsilon_{i}{\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})$
$\Longrightarrow\ \upsilon_{i}\leq\dfrac{2(\epsilon_{L}-1){\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})}{{\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}}.$
where positive ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ is considered. Assuming the Lipschitz gradient continuity of $L({\bm{\rho}})$, $\|\nabla^{2}L({\bm{\rho}}_{i})\|$ is upper bounded. This indicates a positive and finite ${\bm{d}}_{i}^{\rm T}\nabla^{2}L({\bm{\rho}}_{i}){\bm{d}}_{i}$ and thus non-zero $\upsilon_{i}$. Furthermore, in light of (44), we have
$\small\Delta^{2}L({\bm{\rho}}_{i+1})\leq\epsilon_{L}^{2}\upsilon_{i}^{2}({\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i}))^{2}.$ (65)
Since $L({\bm{\rho}})$ is a monotonously decreasing and bounded function regarding ${\bm{\rho}}_{i}$, the (conjugate) gradient-based descent algorithm converges to $\Delta^{2}L({\bm{\rho}}_{i+1})\rightarrow 0$ when $i>I$ for a large constant $I$. However, the deflection factor computed by (41) or the descent direction vector in (49) can guarantee a strict descent ${\bm{d}}_{i}^{\rm T}\nabla L({\bm{\rho}}_{i})<0$ at each iteration $i$ unless ${\bm{d}}_{i}={\bm{0}}$. Consequently, sequence $\{L({\bm{\rho}}_{i})\}$ will converge to a stationary point with a zero descent direction ${\bm{d}}_{i}\rightarrow{\bm{0}}$ when it has Lipschitz gradient continuity. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ when ${\bm{d}}_{i}\rightarrow{\bm{0}}$. On the other hand, we assume the objective function is not Lipschitz gradient continuous. In this case, if the sequence $\{{\bm{\rho}}_{i}\}$ pass a Lipschitz noncontinuous point ${\bm{\rho}}_{I}$ where ${\bm{d}}_{I}^{\rm T}\nabla^{2}L({\bm{\rho}}_{I}){\bm{d}}_{I}$ is extremely large, to satisfy (C), the step size should meet $\upsilon_{I}\rightarrow 0$. According to Lemma 1, we have $\|\Delta{\bm{\rho}}_{i+1}\|\rightarrow 0$ and $\Delta L({\bm{\rho}}_{i+1})\rightarrow 0$ when $\upsilon_{i}\rightarrow 0$ for $i\geq I$, i.e., the sequences $\{{\bm{\rho}}_{i}\}$ and $\{L({\bm{\rho}}_{i})\}$ converge to a Lipschitz gradient noncontinuous point. $\blacksquare$